Graph the following system of equations. Describe both lines and the final graph. (example: The first line is dotted, falls left to right and has shading above it.)
x + y < -2 and x - y < 4

Answer by lwsshak3(6490)

Graph the following system of equations. Describe both lines and the final graph. (example: The first line is dotted, falls left to right and has shading above it.)
x + y < -2 and x - y < 4
**
convert inequalities to straight line equations.
x+y=-2
y=-x-2
This is a straight line with a slope=-1 and y-intercept=-2
For the inequality, shade the area below the line but not including the line(dotted)
..
x-y=4
y=x-4
This is a straight line with a slope=1 and y-intercept=-4
For the inequality, shade the area below the line but not including the line(dotted)